OpenStudy (anonymous):

Which type of triangle has side lengths of 5 in, 6 in, and 7 in?

OpenStudy (anonymous):

Well do you know the difference between isosceles, scalene, and equilateral triangles?

OpenStudy (anonymous):

yes ..

OpenStudy (anonymous):

Well with the sides it says what type of triangle do you think it is :)?

OpenStudy (anonymous):

Remember it has 3 different side lengths

OpenStudy (anonymous):

my choices are , acute , equilateral , obtuse , or right ? so whats ya answer ?

OpenStudy (anonymous):

Well can we try and eliminate some? Like which one would it definitely not be?

OpenStudy (anonymous):

equilateral .. ? obtuse ?

OpenStudy (anonymous):

equilateral we can definitely eliminate.

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

is it a acute ?

OpenStudy (anonymous):

What we can use is a^2+b^2=c^2 which is 5^2+6^2= can you find that?

OpenStudy (anonymous):

And then after you find that you'll take the third side (7^2 or =c^2). If the answer you get is less than C it is obtuse. If it is more it is acute. And if it is the same then it is a right triangle :)